6. Determine the sets on which the following functions are analytic. a. f(z) = Log(z + 1) b. f(x) = Log(x+)
Answer #1


Given f (z)= Log(z+1) Logz isnot analytic at the points z = x+iy, where xs 0 and y = 0. Log(2+1) = Log(1+x+i(y))is not analyt